OpenStudy (anonymous):

determine whether the followin function is linear transformation . T:f(negative infinty,infinity)->f(negative infinity,infinity) where T(f(x))=1+f(x)

OpenStudy (anonymous):

I think it must be this T(f(x)),1+f(x)). Isn't it?

OpenStudy (anonymous):

T(f(x))=1+f(x)

OpenStudy (turingtest):

two criteria to check...

OpenStudy (turingtest):

\[T(c\vec u)=cT(\vec u)\]and\[T(\vec u+\vec v)=T(\vec u)+T(\vec v)\]

OpenStudy (turingtest):

if both these things are true it's a linear transformation, on at least one is not, it aint

OpenStudy (turingtest):

what does the transformation to to the function?

OpenStudy (turingtest):

...in this case I mean

OpenStudy (anonymous):

according to my understanding,is a structure preserving map called hormomophism

OpenStudy (turingtest):

yeah, but those names won't help you if you don't understand how to check them

OpenStudy (turingtest):

what does T do to the function f(x) ???

OpenStudy (anonymous):

transform a function

OpenStudy (turingtest):

how exactly does *this* transformation alter the function?

OpenStudy (anonymous):

it alters a function by the above two conditions u mentioned,ie if it meets the two conditions it a a linear transformation which transforms a function

OpenStudy (turingtest):

you are not understanding, and hence over-complicating my question

OpenStudy (turingtest):

you are told that the transformation you have is\[T(f(x))=f(x)+1\]so this transformation takes a vector (in this case a function) and adds 1 to it

OpenStudy (turingtest):

so given that the operation of this transformation is to add 1 to the input, what is\[T(cf(x))\]where c is some constant

OpenStudy (anonymous):

yah,continue maybe i will understand

OpenStudy (turingtest):

look, what is the input of the transform I just wrote above?

OpenStudy (anonymous):

negative infinity and infinity

OpenStudy (turingtest):

no, it's the vector; the part in the parentheses in the T...

OpenStudy (turingtest):

for\[T(f(x))=f(x)+1\]the input is f(x) and the transformation T is to add 1 to it

OpenStudy (turingtest):

so what is the input in\[T(cf(x))\]

OpenStudy (anonymous):

cf(x)

OpenStudy (turingtest):

right, now apply the transformation and what do you get?

OpenStudy (anonymous):

cf(x)+1

OpenStudy (turingtest):

exactly, so now we know that\[T(cf(x))=cf(x)+1\]now we need to see if that is the same as\[cT(f(x))\]can you check that?

OpenStudy (anonymous):

is nt the same since it gives c(f(x)+1)=cf(x)+c

OpenStudy (turingtest):

exactly :) the only c that would make this linear is c=1, which we cannot assume, so the transformation is not linear you could also have seen that if you tried the other test as well

OpenStudy (anonymous):

lets try the 2nd condition

OpenStudy (turingtest):

ok, what is\[T(f(x)+g(x))\]?

OpenStudy (anonymous):

(f(x)+g(x))+1

OpenStudy (turingtest):

exactly, and what is\[T(f(x))+T(g(x))\]?

OpenStudy (anonymous):

(f(x)+1)+(g(x)+1)

OpenStudy (turingtest):

which simplifies to...?

OpenStudy (anonymous):

f(x)+g(x)+2

OpenStudy (turingtest):

so is\[T(f(x)+g(x))=T(f(x))+T(g(x))\]?

OpenStudy (anonymous):

no

OpenStudy (turingtest):

so it does not fit either condition for a linear transformation, though as I said we were done once we say it did not meet at least one criteria

OpenStudy (anonymous):

thanks a bunch

OpenStudy (turingtest):

welcome!